whwt is the solution to this sequence
1 8 27 64 .....

Answer by stanbon(75887)

whwt is the solution to this sequence
1 8 27 64 .....
------
1^3 = 1
2^3 = 8
3^3 = 27
4^3 = 64
==============
Cheers,
Stan H.
==============